OpenStudy (madmerc):

Which of the following represents the zeros of f(x) = x3 − 4x2 − 5x + 20?

OpenStudy (danjs):

f(x) = 0 x-axis intercepts

OpenStudy (danjs):

looks like one of those where you can regroup things, and factor it

OpenStudy (madmerc):

the choices are −4, −square root of 5, −square root of 5 4, square root of 5, −square root of 5 4, −square root of 5, −square root of 5 −4, square root of 5, square root of 5

OpenStudy (madmerc):

i don't understand how to get to those

OpenStudy (danjs):

[x^3 - 5x] - 4x^2 + 20 both of those pairs have a common factor x*(x^2 - 5) - 4 * (x^2 - 5)

OpenStudy (danjs):

so you can then factor that now, take x^2 -5 from both = (x^2 - 5) * ( x - 4 )

OpenStudy (madmerc):

what do I do after that? @DanJS

OpenStudy (aihberkhan):

That is it.

OpenStudy (aihberkhan):

The answer is: \[(x^2-5)(x-4)\] and @DanJS Is now offline.

OpenStudy (aihberkhan):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @madmerc

OpenStudy (madmerc):

@AihberKhan that isn't one of the choices though..

OpenStudy (madmerc):

The answer was 4, square root of 5, −square root of 5

OpenStudy (mathmate):

hint: use identity: a^2-b^2=(a+b)(a-b). So if \((x^2-7)=0\), we have \((x^2-(\sqrt 7)^2)=(x+\sqrt 7)(x-\sqrt 7)=0 ~~~=>~~~ x=\sqrt 7~or~ x=-\sqrt 7\)